Find the characteristic polynomial of A=
Answer #1

A=[..3 .: 2X2 14 2 characteristic polynomian P (2)=det (A-212) so, p(a) 11-0: 3 1 . 14:22 =(1-) (2-1) -12 = 2-2-2 À -12 =-39-